OpenStudy (anonymous):

9r^2+15r+6 factored out with work

OpenStudy (anonymous):

would the solution (9r+3)(r+2) be correct

OpenStudy (dumbcow):

not quite 18r+3r = 21r not 15r

OpenStudy (anonymous):

Let's work it out. (9r + 3)(r + 2) First: \[9r*r = 9r ^{2}\] Outside: 9r*2=18r Inside: 3*r = 3r Last: 3*2=6 \[9r ^{2}+18r + 3r + 6 = 9r^{2} + 21r + 6\]

OpenStudy (anonymous):

So, not quite. As dumbcow beat me to it. :)

OpenStudy (dumbcow):

9*6=54 9+6=15 (9r+6)(9r+9) reduce (3r+2)(3r+3)

OpenStudy (anonymous):

ohhhh

OpenStudy (dumbcow):

actually i forgot you could factor out a 3 first 3(3r^2 +5r +2) =3(3r+2)(r+1)

OpenStudy (anonymous):

ok lol